(a) What is the average translational kinetic energy of an ideal-gas molecule at 27°C?
(b) What is the total random translational kinetic energy of the molecules in 1 mole of this gas?
(c) What is the root-mean-square speed of oxygen molecules at this temperature?
1
Expert's answer
2019-05-03T09:07:22-0400

(a) The average translational kinetic energy of an ideal-gas molecule is calculated as


"E=(3\/2)kT"

where T is a temperature in Kelvins

"Tk \approx Tc+273"

, and

"k=1.38\u00d710^{\u221223} J\/K"

is a Boltzmann constant

So we can calculate the energy as

"E=(3\/2)kT=(3\/2)*1.38\u00d710^(\u221223)*(273.3+27)"

"E=6.21*10^{e-21}"

(b) The total random translational kinetic energy of the molecules in 1 mole of this gas is equal to the sum of the energies of all molecules,

"E_{total}=\sum{E}=E*N_{Avogadro}"

"E=6,023*(10^23)*6.21*10^(-21)=3740283J"

where Navogadro is the number of molecules in one mole (Navogadro= 6,023*10^(23))

(c) The root-mean-square speed of oxygen molecules at this temperature can be calculated from the relation

"E_{trans}=mv^2\/2=(3\/2)kT"

where m is the mass of the oxygen molecule (m(O-molecule) = 2*m(O-atom) = 2.66*10^(-26))

"v= \sqrt{3kT/m}"

"v= \sqrt{3*1.38*10^{-23}*300/{(2.66*10^{-26})}}=683.31m/s"